OpenStudy (anonymous):

180 miles/2 hours find the unit rate in feet per second

OpenStudy (anonymous):

well you need to set up a proprotion

OpenStudy (jdoe0001):

\(\bf \cfrac{180\cancel{mi}}{2\cancel{hr}}\cdot \cfrac{5280ft}{\cancel{mi}}\cdot \cfrac{\cancel{hr}}{60\cancel{min}}\cdot \cfrac{\cancel{min}}{60secs}\)

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

Thanks man have a good day

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[\frac{180*5280}{2*3600}\text{=}132 \text{ ft}/\sec\]